OpenStudy (anonymous):

Express the equation in slope-intercept form: 7x-5y=10

OpenStudy (aggie12341):

y=5/7x-2

OpenStudy (anonymous):

i dont have that as an anwser choice

OpenStudy (anonymous):

3x-x+2=4

OpenStudy (aggie12341):

y=-7/5x-2

OpenStudy (anonymous):

that either i have 7/5x+-2

OpenStudy (aggie12341):

that's fine

OpenStudy (aggie12341):

is there a -7/5

OpenStudy (anonymous):

Slope-intercept form: y = mx + c 7x-5y=10 -5y=10-7x Divide it by -5 to get y alone: y = 7x/5 + 2

OpenStudy (anonymous):

ok

OpenStudy (aggie12341):

wrong @Cupcakescankill the variable xgoes first

OpenStudy (anonymous):

Sorry my bad its y = (7 x)/5-2